# SMART: Simulated Students Aligned with Item Response Theory for Question Difficulty Prediction

### v1

@@ A.1 Student Simulation Hyperparameters
We perform a preliminary hyperparameter search on our validation sets, and use the same hyperparameters on all models for consistency. We set LoRA’s rank = 128, $\alpha$ = 64, and dropout = 0.05. We perform SFT for 1 epoch, with learning rate = 1e-4 and linear warmup for 10% of steps, weight decay = 1e-2, gradient clipping norm = 1.0, and effective batch size = 64 using gradient accumulation. For DPO Rafailov et al. (2023), we set learning rate = 1e-5, $\beta$ = 0.5, the probability threshold $\epsilon=0.1$ (Equation 3), the number of negative candidates $m=3$, and train on a random 20% subset of the training set to reduce costs. We conduct all experiments on 8 NVIDIA A10G GPUs.
At inference time, we sample 1,000 student abilities for simulation and generate a single response for each item/ability pair. We decode up to 500 tokens with nucleus sampling, setting temperature = $0.7$ and p = $0.95$. We use an inference-time batch size of 64 for the simulated student model and 2 for the scoring model, optimizing for inference speed.
For both SFT and DPO, training the simulated student model takes approximately 3 hours on a single GPU for the 1B model and 6 hours on 2 GPUs for the 3B model. Training the automated scoring model takes approximately 4 hours on a single GPU for the 1B model and 9 hours on 2 GPUs for the 3B model. Generating 10,000 responses (1,000 abilities over 10 items) takes approximately 1 hour on a single GPU for the 1B model and 3 hours on a single GPU for the 3B model. Scoring these responses takes approximately 15 minutes on a single GPU for the 1B model and 35 minutes on a single GPU for the 3B model.
@@ A.2 Additional Implementation Details
For our Response Gen Zero-shot student simulation baseline, we first normalize all train abilities to $[-3,3]$ before sampling, and then map abilities to 10 labeled buckets with the following ranges: “Minimal”: $\theta<-2.5$, “Emerging”: $\theta\in[-2.5,-2)$, “Developing”: $\theta\in[-2,-1.5)$, “Approaching Proficiency”: $\theta\in[-1.5,-0.8)$, “Proficient”: $\theta\in[-0.8,0)$, “Competent”: $\theta\in[0,0.8)$, “Skilled”: $\theta\in[0.8,1.5)$, “Advanced”: $\theta\in[1.5,2.0)$, “Exceptional”: $\theta\in[2,2.5)$, “Mastery”: $\theta\geq 2.5$.
@@ A.3 Predicted Difficulty Normalization
$\hat{b}_{i}=\frac{\sigma_{2}}{\sigma_{1}}(\hat{b}_{i}-\mu_{1})+\mu_{2},$ (4)
@@ A.4 IRT Model Fitting
We implement our GPCM-based IRT model using PyTorch, using the formulation in Equation 2 to compute the probability of a scored response. We do not enforce normal priors on any parameters, and use cross-entropy loss with the ground truth scores and gradient descent to train the model. We train using the AdamW optimizer for 50 epochs with learning rate = 1e-3, weight decay = 0, and batch size = 256. We reserve a random set of 20% of all responses as a validation set, and compute the quadratic weighted kappa (QWK) between the most likely scores predicted by the IRT model and the ground truth scores in the validation set. The average QWK across the test sets of all folds is $0.57$, showing moderate-to-high agreement.
@@
Our dataset does not contain student identifiers, but rather contains prior ability estimates based on each student’s responses to a larger set of questions, including other types such as multiple choice questions; the 49 short answer questions we use comprise all short answer items in the full set. To ensure that other item types and test set responses do not influence our ability estimates, we fit an IRT model from scratch only on the 49 items used in this study. Since we do not have access to student identifiers, we use the prior ability as a proxy. However, the student responses are very sparse on our set of questions, with around $63$K unique ability values for around $85$K responses, leading to a poor fit for the IRT model. As a solution, we round each prior ability to a single decimal place (we found this precision to work best out of several tried) and use the resulting value as the student identifier. This transformation leads to a significantly better fit for the IRT model. We note that this process assumes access to prior estimates of student ability, which may or may not be available in real-world settings. However, we note that this a limitation of the data, rather than our method. Additionally, we note that this transformation does not unfairly benefit our method over the baselines, so the comparison between methods in this work remains fair.
@@ B.1 Cross-Validation Splitting
We perform our cross-validation splits such that 1) each item appears in train, validation, and test at least once, and 2) the distribution of item difficulties should be roughly the same across train, validation, and test over all folds. We achieve these requirements using the following algorithm. First, we sort all items according to their difficulty. Second, we partition all items into 10 difficulty buckets of up to 5 items each. Third, we sequentially construct a list of items by adding one random item at a time from each bucket in a round-robin fashion; the resulting list is striped by difficulty bucket. Finally, we evenly rotate this list 5 times to form the 5 cross-validation folds, partitioning the list at 29 and 39 items to create the train/validation/test splits.
@@ B.2 Passage Summarization
Several questions include long reading passages, up to $4,000$ tokens in length, leading to increased memory usage and training time; therefore, we prompt Llama-3.1-70B-Instruct, loaded with NF4 quantization, to summarize these passages while retaining the main elements and original writing style. We show this prompt in Appendix E.
Appendix C Statement on Scientific Artifacts
@@
To the best of our knowledge, all software and models we build our implementation on have open-source licenses or no available license. Additionally, we are within their intended terms of use, specifically, scientific research. If we release code, we will ensure the license and terms of use reflect the sources we build on.
@@
Our dataset is a proprietary and contains private student data belonging to a state-level consortia. We do not examine the data for offensive content or identifiable information, because 1) the data is private and securely stored, 2) all models trained on the data are securely stored and will not be released, and 3) the data is very large and would be difficult to effectively scan for such issues.
Appendix D Qualitative Examples
@@
Middle In source 1 it talks about a robot named "Gemini Scout" that is a remote controlled robot. It can also be used for mine, farms, and doctors. It can also help people who are in danger. It can also help with rescues. In source 2 it talks about…
@@
Middle In source 1 it talks about a robot named "Gemini Scout" that can carry food and water. It can also collect information about the air that comes out of the vents in the volcano. In source 2 it talks about…
@@
Table 2: Examples of student responses generated by SFT and SMART for a public item, conditioned on abilities in the low, middle, and high ranges. SMART adapts better to the instructed ability than SFT. For low ability, SMART introduces a spelling error and poor capitalization, and does not back up its claim, while SFT does not answer the question correctly but uses better sentence structure. For medium ability, both methods provide evidence but do not explain how they support the claim. For high ability, SMART produces structured writing, provides evidence, and explains how its evidence supports the claim, while SFT is less explicit about how its evidence supports the claim.
@@
Middle Plants are very important to us because they give us oxygen, that we need to survive. Plants are very important because they give us oxygen.
@@
Middle Plants are one of the most important things on earth, without them we would not be alive today. Plants are very important to us humans because without them we wouldn’t be able to survive, they also help us in many ways like we can make food out of them and plants help us by providing oxygen.
@@
Table 3: Examples of student responses generated by SFT and SMART for a public item, conditioned on abilities in the low, middle, and high ranges. SMART adapts better to the instructed ability than SFT. For low ability, SMART uses very simple language and doesn’t discuss how plants make food, while SFT uses more formal language and does mention how plans make food, which would be atypical for a low ability student. For medium ability, both methods write a coherent introduction but don’t mention how plants make food; SMART writes a slightly longer response. For high ability, SMART makes sure to motivate the importance of photosynthesis as well as directly answer the question by stating how plants make food, while SFT neglects to mention how plants make food.
Appendix E Prompts

### v2

@@ A.1 Student Simulation Hyperparameters
We perform a preliminary hyperparameter search on our validation sets. We set LoRA’s rank = 128, $\alpha$ = 64, and dropout = 0.05. For Smarter Balanced, for both the scoring model and student model, we perform SFT for 1 epoch, with learning rate = 1e-4 and linear warmup for 10% of steps, weight decay = 1e-2, gradient clipping norm = 1.0, and effective batch size = 64 using gradient accumulation. For DPO, we set learning rate = 1e-5, $\beta$ = 0.5, the probability threshold $\epsilon=0.1$ (Eq. 4), the number of negative candidates $m=3$, and train on a random 20% subset of the training set to reduce costs. For CodeWorkout, we use the same hyperparamters except for the following: for the scoring model we perform SFT for 5 epochs with learning rate = 5e-5, for the student model we perform SFT for 3 epochs with learning rate = 1e-5, and perform DPO for 1 epoch with learning rate = 1e-6 on the full training set. We conduct all experiments on the Smarter Balanced dataset on NVIDIA A10G GPUs and all experiments on the CodeWorkout dataset on NVIDIA A40 GPUs.
At inference time, we sample $n=1,000$ student abilities for simulation and generate a single response for each item/ability pair. We decode up to $500$ tokens with nucleus sampling, setting temperature = $0.7$ and p = $0.95$. We use an inference-time batch size of $64$ for the simulated student model and $2$ for the scoring model, optimizing for inference speed.
On the Smarter Balanced dataset, for both SFT and DPO, training the simulated student model takes approximately 3 hours on a single GPU for the 1B model and 6 hours on 2 GPUs for the 3B model. Training the automated scoring model takes approximately 4 hours on a single GPU for the 1B model and 9 hours on 2 GPUs for the 3B model. Generating 10,000 responses (1,000 abilities over 10 items) takes approximately 1 hour on a single GPU for the 1B model and 3 hours on a single GPU for the 3B model. Scoring these responses takes approximately 15 minutes on a single GPU for the 1B model and 35 minutes on a single GPU for the 3B model.
@@ A.2 Additional Implementation Details
For our Response Gen ZS student simulation baseline, we first normalize all train abilities to [3,3][-3,3] before sampling, and then map abilities to 10 labeled buckets with the following ranges: “Minimal”: $\theta<-2.5$, “Emerging”: $\theta\in[-2.5,-2)$, “Developing”: $\theta\in[-2,-1.5)$, “Approaching Proficiency”: $\theta\in[-1.5,-0.8)$, “Proficient”: $\theta\in[-0.8,0)$, “Competent”: $\theta\in[0,0.8)$, “Skilled”: $\theta\in[0.8,1.5)$, “Advanced”: $\theta\in[1.5,2.0)$, “Exceptional”: $\theta\in[2,2.5)$, “Mastery”: $\theta\geq 2.5$.
@@ A.3 Predicted Difficulty Normalization
$\hat{b}_{i}\leftarrow\frac{\sigma_{2}}{\sigma_{1}}(\hat{b}_{i}-\mu_{1})+\mu_{2},$ (6)
@@ A.4 IRT Model Fitting
The Smarter Balanced dataset does not contain student identifiers, but rather contains prior ability estimates based on each student’s responses to a larger set of questions, including other types such as multiple choice questions; the 49 short answer questions we use comprise all short answer items in the full set. To ensure that other item types and test set responses do not influence our ability estimates, we fit an IRT model from scratch only on the 49 items used in this study. Since we do not have access to student identifiers, we use the prior ability as a proxy. However, the student responses are very sparse on our set of questions, with around 6363K unique ability values for around 8585K responses, leading to a poor fit for the IRT model. As a solution, we round each prior ability to a single decimal place (we found this precision to work best out of several tried) and use the resulting value as the student identifier. This transformation leads to a significantly better fit for the ground-truth IRT model. We note that this process assumes access to prior estimates of student ability, which may or may not be available in real-world settings. However, we note that this a limitation of the data, rather than our method. Additionally, we note that this transformation does not unfairly benefit our method over the baselines, so the comparison between methods in this work remains fair.
@@
We implement our IRT models using PyTorch, using the formulations in Equations 2 and 3 to compute the probability of a student’s score for a particular item. For GPCM, following prior work Muraki (1993), we set the first step parameter to $d_{i0}=0$, since only C1C-1 step parameters are required for $C$ score classes. We also impose the location constraint Muraki (1993), where all step parameters for an item sum to 0, i.e., $\sum_{y=0}^{C-1}d_{iy}=0$, ensuring that difficulty parameters across items are on the same scale and are identifiable. We also fix all discrimination parameters to $1$, which improves model fit. We do not enforce normal priors on any parameters. We train the model using gradient descent, using the cross-entropy loss between the predicted score probabilities and the ground-truth scores. We train using the AdamW optimizer for 50 epochs with learning rate = 1e-3, weight decay = 0, and batch size = 256; for CodeWorkout, we train for 200 epochs when fitting the ground-truth data, which is necessary due to the smaller number of ground-truth student responses. We reserve a random validation set of 20% of all responses to determine how predictive our ground-truth IRT parameters are, comparing the most likely student scores predicted by the IRT model and the ground-truth scores in the validation set. For Smarter Balanced, we measure performance using QWK since scores are ordinal, with the IRT model achieving a QWK of $0.57$ on the validation set. For CodeWorkout, we measure performance using PCC since scores are continuous, with the IRT model achieving a PCC of $0.49$. These values show moderate predictability, indicating that while the student scores are difficult to predict, the parameters are well-calibrated.
@@ B.1 Cross-Validation Splitting
We create our cross-validation splits such that 1) each item appears in train, validation, and test at least once, and 2) the distribution of item difficulties should be roughly the same across train, validation, and test over all folds. We achieve these requirements using the following algorithm. First, we sort all items according to their difficulty. Second, we partition all items into 10 difficulty buckets of up to 5 items each. Third, we sequentially construct a list of items by adding one random item at a time from each bucket in a round-robin fashion; the resulting list is striped by difficulty bucket. Finally, we evenly rotate this list 5 times to form the 5 cross-validation folds, partitioning each rotation at 60% and 80% to create a train/validation/test split.
@@ B.2 Passage Summarization
Several Smarter Balanced questions include long reading passages, up to $4,000$ tokens in length, leading to increased memory usage and training time; therefore, we prompt Llama-3.1-70B-Instruct, loaded with NF4 quantization, to summarize these passages while retaining the main elements and original writing style. We show this prompt in Appendix F.
Appendix C Data Scaling Experiment
Figure 3: Results of data scaling experiment. Even a small amount of data is sufficient to achieve high performance. SMART begins to outperform SFT more significantly when more data is available.
We conduct an experiment where we vary the amount of data available for training by using random subsets of the student responses in the training set at 0%, 25%, 50%, 75%, and 100%. We re-train both SFT and SMART using the 1B model on these subsets, and report the PCC averaged over all cross-validation test splits on CodeWorkout. To only examine the impact of training data on the student model, we use the scoring model trained on all data, since in reality a stronger scoring model may be available (such as one that runs test cases on generated code).
We show the results of this data scaling experiment in Figure 3. We observe that even a small amount of training data is sufficient to drastically increase performance from zero-shot (25% of the training data contains 1,636 responses and 3,027 preference pairs in total). SFT and SMART perform similarly until 75%, when SMART begins to outperform SFT. Surprisingly, performance drops for both methods from 75% to 100%, indicating that some training responses may negatively impact performance, such as ones that represent item types that are not present in the test set. This result further reflects the challenges of testing on a small number of items.
Appendix D Statement on Scientific Artifacts
@@
To the best of our knowledge, all software and models we build our implementation on have open-source licenses or no available license. Additionally, we are within their intended terms of use, specifically, scientific research. If we release code, we will ensure the license and terms reflect the sources we build on.
@@
The Smarter Balanced dataset contains proprietary questions and private student data. We do not examine the data for offensive content or identifiable information, because 1) the data is private and securely stored, 2) all models trained on the data are securely stored and will not be released, and 3) the data is very large and would be difficult to effectively scan for such issues. We do not check for such issues on CodeWorkout either, because it is a public and widely used dataset, and such investigations are beyond the scope of this work.
Appendix E Qualitative Examples
@@
Medium In source 1 it talks about a robot named "Gemini Scout" that is a remote controlled robot. It can also be used for mine, farms, and doctors. It can also help people who are in danger. It can also help with rescues. In source 2 it talks about…
@@
Medium In source 1 it talks about a robot named "Gemini Scout" that can carry food and water. It can also collect information about the air that comes out of the vents in the volcano. In source 2 it talks about…
@@
Table 3: Examples of student responses generated by SFT and SMART for a publicly available Smarter Balanced item, conditioned on abilities in the low, medium, and high ranges. SMART adapts better to the instructed ability than SFT. For low ability, SMART introduces a spelling error and poor capitalization, and does not back up its claim, while SFT does not answer the question correctly but uses better sentence structure. For medium ability, both methods provide evidence but do not explain how they support the claim. For high ability, SMART produces structured writing, provides evidence, and explains how its evidence supports the claim, while SFT is less explicit about how its evidence supports the claim.
@@
Medium Plants are very important to us because they give us oxygen, that we need to survive. Plants are very important because they give us oxygen.
@@
Medium Plants are one of the most important things on earth, without them we would not be alive today. Plants are very important to us humans because without them we wouldn’t be able to survive, they also help us in many ways like we can make food out of them and plants help us by providing oxygen.
@@
Table 4: Examples of student responses generated by SFT and SMART for a publicly available Smarter Balanced item, conditioned on abilities in the low, medium, and high ranges. SMART adapts better to the instructed ability than SFT. For low ability, SMART uses very simple language and doesn’t discuss how plants make food, while SFT uses more formal language and does mention how plans make food, which would be atypical for a low ability student. For medium ability, both methods write a coherent introduction but don’t mention how plants make food; SMART writes a slightly longer response. For high ability, SMART makes sure to motivate the importance of photosynthesis as well as directly answer the question by stating how plants make food, while SFT neglects to mention how plants make food.
Table 5: Examples of real student responses and responses generated by SFT and SMART for a CodeWorkout item, conditioned on abilities in the low, medium, and high ranges. Scores for real students represent the portion of test cases that passed; scores for simulated students are assigned by our trained scoring model. The true difficulty for the question is -0.57, the SFT difficulty prediction is 0.22, and the SMART difficulty prediction is -0.41. SMART responses are much more diverse than SFT responses, likely resulting in the more accurate prediction. To further demonstrate this, we report the average score for each ability quartile from each source, where real students and SMART show a gradual increase in score with ability, while SFT does not show any distinction between abilities. Real students: [0.22, 0.53, 0.71, 0.95], SFT: [0.81, 0.79, 0.82, 0.83], SMART: [0.57, 0.79, 0.85, 0.93].
Question Return the centered average of an array of ints, which we’ll say is the mean average of the values, except ignoring the largest and smallest values in the array. If there are multiple copies of the smallest value, ignore just one copy, and likewise for the largest value. Use int division to produce the final average. You may assume that the array has a length of 3 or more.
Ability Response Score
Human Student
Low
public int centeredAverage(int[] nums)
    int merlin = 0;
    int less = nums[0];
    int more = nums[0];
    for (int i = 0; i < nums.length; i++)
        merlin = merlin + nums[i];
        less = Math.min(less, nums[i]);
        more = Math.max(more, nums[i]);
    return (sum - (min-max)/nums.length);
0.0
Medium
public int centeredAverage(int[] nums)
    int max = 0;
    int min = nums[0];
    for (int i = 0; i < nums.length; i++)
        if (nums[i] > max)
            max = nums[i];
        else if (nums[i] < min)
            min = nums[i];
    int sum = 0;
    int ct = 0;
    for (int n : nums)
        if (n != min && n != max)
            sum += n;
            ct++;
    return sum/ct;
0.67
High
public int centeredAverage(int[] nums)
    int sum = 0;
    int min = nums[0];
    int max = nums[0];
    for(int i = 0; i < nums.length; i++) {
        sum += nums[i];
        min = Math.min(min, nums[i]);
        max = Math.max(max, nums[i]);
    return (sum - min - max) / (nums.length - 2);
1.0
SFT
Low
public int centeredAverage(int[] nums)
    int sum = 0;
    for (int i = 0; i < nums.length - 2; i++)
        sum = sum + nums[i];
    return sum / (nums.length - 2);
0.68
Medium
public int centeredAverage(int[] nums)
    int sum = 0;
    for (int i = 0; i < nums.length; i++)
        sum = sum + nums[i];
    int avg = sum / (nums.length - 2);
    return avg;
0.75
High
public int centeredAverage(int[] nums)
    int sum = 0;
    for (int i = 0; i < nums.length; i++) {
        if (nums[i]!= Integer.MIN_VALUE && nums[i]!= Integer.MAX_VALUE)
            sum += nums[i];
    return sum / (nums.length - 2);
0.73
SMART (ours)
Low
public int centeredAverage(int[] nums)
    int i = 0;
    int sum = 0;
    for (int i = 0; i < nums.length; i++)
        sum = sum + nums[i];
    return sum / nums.length - 1;
0.27
Medium
public int centeredAverage(int[] nums)
    int sum = 0;
    int min = Integer.MAX_VALUE;
    int max = Integer.MIN_VALUE;
    for (int i = 0; i < nums.length; i++) {
        if (nums[i] > max) {
            max = nums[i];
        if (nums[i] < min) {
            min = nums[i];
        sum += nums[i];
    return (sum - max - min) / (nums.length - 2);
0.94
High
public int centeredAverage(int[] nums) {
  int min = Integer.MAX_VALUE;
  int max = Integer.MIN_VALUE;
  int sum = 0;
  for(int num : nums) {
    if(num < min) min = num;
    if(num > max) max = num;
    sum += num;
  sum -= min + max;
  return sum / (nums.length - 2);
0.98
Appendix F Prompts